40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 16 02:11:40 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged 
    Reply  Wed Apr 16 02:29:30 2014, ericq, Update, LSC, CARM and DARM on IR signals, boosts engaged carmOLTF.pdfboostPlot.png
       Reply  Thu Apr 17 00:49:06 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged Lockloss_DARMgainTooHigh_119am.pngLockloss_DARMgainTooHigh_127am.pngSmoothCARMandDARMtransitions_LSCboosts.png
Message ID: 9818     Entry time: Wed Apr 16 02:29:30 2014     In reply to: 9817     Reply to this: 9819
Author: ericq 
Type: Update 
Category: LSC 
Subject: CARM and DARM on IR signals, boosts engaged 

 As Jenne mentioned, we took OLTF transfer functions, and determined that we had more than enough phase margin to switch on the LSC boosts in FM4. This improved the error signal noise spectra quite a lot, and noticeably reduced the TRX/TRY fluctuations, and actuation output. 

Here's the CARM OLTF (FM4 boost on in red, boost off in black)

carmOLTF.pdf

 

Here's what happened to the CARM and DARM spectra when we turned on the boosts. (ALS only in black, initial IR signal transitions in mid-color, boosted IR signals in bright color)

boostPlot.png 

ELOG V3.1.3-